# Do deep neural networks behave like structural glasses?

### v3

@@ I Introduction
The goal of this work is to quantitatively assess the similarities between DNNs and structural glass-forming materials by performing comparisons across a wide range of numerical experiments typical of glassy formers. Do DNNs undergo glass-like transitions? Do they exhibit the typical glassy behaviours of diverging relaxation times, caging, violation of the Stokes-Einstein relation, dynamic heterogeneity, and ageing? Answering these questions is an important first step in guiding the application of the theory of structural glasses to the field of deep learning.
@@ III.1 Phase diagram
In order to obtain the phase diagram, we sample a uniform grid in the $(N,\lambda)$ plane, i.e. construct DNNs of varying size and regularisation, and train these networks with Stochastic Gradient Descent for 7000 epochs with a batch size of 100. We then perform 50 epochs of gradient descent at the original learning rate, and a further 50 epochs of gradient descent after reducing the learning rate by a factor of 10. Convergence was checked manually on a representative subset of DNNs. As such we are sampling (close to) the minima of the landscape. Finally we average over 10 DNNs with different, random initial conditions at each point in the $(N,\lambda)$ plane, resulting in the phase diagram shown in Fig. 1(a). We use SGD instead of Langevin dynamics as we are studying static properties of the landscape, not the dynamics, and SGD is much faster than Eq. (3). We include illustrative sketches of the resulting phases around the edge of Fig. 1(a) emphasising their important properties: the absence/presence of flat regions either side of the UOT, and the emergence of many minima below the TTT.
@@ III.5 The Stokes-Einstein relation and Dynamic Heterogeneity
where the denominator comes from the Stokes drag equation, and the exact constant of proportionality is system dependent [86]. In colloidal glasses (which obey overdamped Langevin dynamics like Eq. (3)) the effective viscosity of the fluid-colloid mixture, $\eta_{\text{eff}}$, is used instead of the bare viscosity of the background fluid [87]. It is a well known signature of the glass transition that the Stokes-Einstein relation begins to be violated as the system enters the supercooled regime (the temperature regime where the system beings to fall out of equilibrium, but has not yet cooled sufficiently to be considered a true glass) [88, 89, 90]. The cause of this violation is thought to be associated with non-Gaussian fluctuations of the particles’ displacement distribution, as measured by the non-Gaussian parameter, $\alpha_{2}$ [91, Kob1997a, Donati1998], and the emergence of dynamic heterogeneity (DH). DH is another typical glassy phenomenon, whereby the system begins to separate into spatially localised populations of mobile and non-mobile particles [90, 92, 93]. The mobile particles dominate the diffusion, whereas the non-mobile particles dominate the viscosity, resulting in the decoupling of these two processes, and the violation of Eq. (9).
@@ IV Discussion
The relationship between deep learning and structural glasses is complicated, with both surprising points of agreement (e.g. DH) and disagreement (e.g. the lack of diverging timescales). Looking beyond structural glasses, the wider field of the physics of disordered systems potentially has a lot to offer as a theoretical blueprint for deep learning. The Random Lorentz Gas and Anderson Localisation both suggest themselves as useful, and as of yet unexplored, physical analogies to DNNs as their complex landscapes result in a localisation transition as opposed to the caging seen in structural glasses. Furthermore, the link between the localisation transition of the RLG and percolation theory is well known, as is the existence of a fractal void space near the percolation transition [106, 107, 108, 109, 110]. Perhaps similar phenomena can be discovered in the connectedness of plateaus in the overparameterised phase. Exploring such topics further would be a highly novel direction of research, and further enrich the exciting work currently being done at the interface of deep learning and statistical physics.
@@ Appendix A Temperature in networks with overdamped Langevin dynamics
We wish to construct the overdamped Langevin dynamics of a network to be as similar as possible to that of a physical system. The degrees of freedom, the weights, are analogous to the particle positions. The energy of a system is an extensive quality, whereas the loss, being an average over the data, is intensive. Consequently, the quantity $P_{D}\mathcal{L}$ plays the role of the energy.
@@ Appendix B Networks above and below the TTT
In Fig. A1 we measure several characteristics of DNNs above and below the TTT. We train 10 networks at $\lambda=0.04$ (high regularisation, above the TTT), and 10 at $\lambda=0.01$ (low regularisation, below the TTT), initialised with weights randomly drawn from a uniform distribution of width $2/\sqrt{N_{L}}$, where $N_{L}$ is the layer width (this is the default PyTorch initialisation). All networks were trained for 7000 epochs using SGD with a learning rate $\mu=10^{-3}$ and batch size 100, followed by 100 epochs of gradient descent to reach their local minima. In Fig. A1(a) we plot the distribution of weights, demonstrating that above the TTT $w_{i}\approx 0$ $\forall i$, whereas below the TTT, the weights adopt a broad range of values. We plot the distribution of final training loss values above the TTT in Fig. A1(b), and below it in Fig. A1(c). These plots demonstrate that above the TTT $\mathcal{L}\approx 0.5$, as expected for the quadratic hinge loss (Eq. (1)) evaluated at $f=0$, whereas below the TTT the loss adopts a spread of different values. Finally, in Fig. A1(d) and (e), we calculate the Euclidean distance in weight space between all pairs of networks, $|\mathbf{W}_{i}-\mathbf{W}_{j}|$, where $\mathbf{W}_{i}$ is the vector of weights of network $i$. Above the TTT (Fig. A1(d)) all networks exist in the $\mathbf{W}=\mathbf{0}$ state, hence the distance between them is zero. Below the TTT, the networks fall into distinct minima, and the distances between them adopt a spread of values. In the main text we use $\mathcal{L}\approx 0.5$ to indicate that the DNNs are in the quadratic state (i.e. above the TTT), as the loss is a convenient and computationally efficient observable. It is highly unlikely, though not impossible, that an ensemble of networks not in the quadratic state, could also all have $\mathcal{L}\approx 0$.

### v4

# Glassy dynamics in deep neural networks: A structural comparison

@@ I Introduction
The goal of this work is to quantitatively assess the similarities between DNNs and structural glass-forming materials by performing comparisons across a wide range of numerical experiments typical of glass formers. Do DNNs undergo glass-like transitions? Do they exhibit the typical glassy behaviours of diverging relaxation times, caging, violation of the Stokes-Einstein relation, dynamic heterogeneity, and ageing? Answering these questions is an important first step in guiding the application of the theory of structural glasses to the field of deep learning.
@@ III.1 Phase diagram
In order to obtain the phase diagram, we sample a uniform grid in the $(N,\lambda)$ plane, i.e. construct DNNs of varying size and regularisation, and train these networks with Stochastic Gradient Descent for 7000 epochs with a batch size of 100. We then perform 50 epochs of gradient descent at the original learning rate, and a further 50 epochs of gradient descent after reducing the learning rate by a factor of 10. Convergence was checked manually on a representative subset of DNNs. As such we are sampling (close to) the minima of the landscape. Finally we average over 10 DNNs with different, random initial conditions at each point in the $(N,\lambda)$ plane, resulting in the phase diagram shown in Fig. 1(a). We use SGD instead of Langevin dynamics as we are studying static properties of the landscape, not the dynamics, and SGD is much faster than Eq. (3). We include illustrative sketches of the resulting phases around the edge of Fig. 1(a) emphasising their important properties: the absence/presence of flat regions on either side of the UOT, and the emergence of many minima below the TTT.
@@ III.5 The Stokes-Einstein relation and Dynamic Heterogeneity
where the denominator comes from the Stokes drag equation, and the exact constant of proportionality is system dependent [86]. In colloidal glasses (which obey overdamped Langevin dynamics like Eq. (3)) the effective viscosity of the fluid-colloid mixture, $\eta_{\text{eff}}$, is used instead of the bare viscosity of the background fluid [87]. It is a well known signature of the glass transition that the Stokes-Einstein relation begins to be violated as the system enters the supercooled regime (the temperature regime where the system begins to fall out of equilibrium, but has not yet cooled sufficiently to be considered a true glass) [88, 89, 90]. The cause of this violation is thought to be associated with non-Gaussian fluctuations of the particles’ displacement distribution, as measured by the non-Gaussian parameter (or excess kurtosis), $\alpha_{2}$ [91, Kob1997a, Donati1998], and the emergence of dynamic heterogeneity (DH). DH is another typical glassy phenomenon, whereby the system begins to separate into spatially localised populations of mobile and non-mobile particles [90, 92, 93]. The mobile particles dominate the diffusion, whereas the non-mobile particles dominate the viscosity, resulting in the decoupling of these two processes, and the violation of Eq. (9).
@@ IV Discussion
The relationship between deep learning and structural glasses is complicated, with both surprising points of agreement (e.g. DH) and disagreement (e.g. the lack of diverging timescales). Looking beyond structural glasses, the wider field of the physics of disordered systems potentially has a lot to offer as a theoretical blueprint for deep learning. The Random Lorentz Gas (RLG) and Anderson Localisation both suggest themselves as useful, and as of yet unexplored, physical analogies to DNNs as their complex landscapes result in a localisation transition as opposed to the caging seen in structural glasses. Furthermore, the link between the localisation transition of the RLG and percolation theory is well known, as is the existence of a fractal void space near the percolation transition [106, 107, 108, 109, 110]. Perhaps similar phenomena can be discovered in the connectedness of plateaus in the overparameterised phase. Exploring such topics further would be a highly novel direction of research, and further enrich the exciting work currently being done at the interface of deep learning and statistical physics.
@@ Appendix A Temperature in networks with overdamped Langevin dynamics
We wish to construct the overdamped Langevin dynamics of a network to be as similar as possible to that of a physical system. The degrees of freedom, the weights, are analogous to the particle positions. The energy of a system is an extensive quantity, whereas the loss, being an average over the data, is intensive. Consequently, the quantity $P_{D}\mathcal{L}$ plays the role of the energy.
@@ Appendix B Networks above and below the TTT
In Fig. A1 we measure several characteristics of DNNs above and below the TTT. We train 10 networks at $\lambda=0.04$ (high regularisation, above the TTT), and 10 at $\lambda=0.01$ (low regularisation, below the TTT), initialised with weights randomly drawn from a uniform distribution of width $2/\sqrt{N_{L}}$, where $N_{L}$ is the layer width (this is the default PyTorch initialisation). All networks were trained for 7000 epochs using SGD with a learning rate $\mu=10^{-3}$ and batch size 100, followed by 100 epochs of gradient descent to reach their local minima. In Fig. A1(a) we plot the distribution of weights, demonstrating that above the TTT $w_{i}\approx 0$ $\forall i$, whereas below the TTT, the weights adopt a broad range of values. We plot the distribution of final training loss values above the TTT in Fig. A1(b), and below it in Fig. A1(c). These plots demonstrate that above the TTT $\mathcal{L}\approx 0.5$, as expected for the quadratic hinge loss (Eq. (1)) evaluated at $f=0$, whereas below the TTT the loss adopts a spread of different values. Finally, in Fig. A1(d) and (e), we calculate the Euclidean distance in weight space between all pairs of networks, $|\mathbf{W}_{i}-\mathbf{W}_{j}|$, where $\mathbf{W}_{i}$ is the vector of weights of network $i$. Above the TTT (Fig. A1(d)) all networks exist in the $\mathbf{W}=\mathbf{0}$ state, hence the distance between them is zero. Below the TTT, the networks fall into distinct minima, and the distances between them adopt a spread of values. In the main text we use $\mathcal{L}\approx 0.5$ to indicate that the DNNs are in the quadratic state (i.e. above the TTT), as the loss is a convenient and computationally efficient observable. It is highly unlikely, though not impossible, that an ensemble of networks not in the quadratic state, could also all have $\mathcal{L}\approx 0.5$.